OpenStudy (anonymous):

The magnitude of the drag force of air resistance on a certain 20.0-kg object is proportional to its speed. If the object has a terminal speed 80.0 m/s, what is the magnitude of the drag force on the object when it is falling with a speed 30.0 m/s?

OpenStudy (anonymous):

Drag force proportional to speed: Fd = kv ; k is some constant We know at terminal velocity Vt: Fg = Fd mg = kVt k = mg / Vt = (20.0)(9.8)/(80.0) = 2.45 kg/s Now we found the constant. We can use that to find the Fd at any speed. Fd = kv = 2.45v Fd = 2.45 (30.0) = 73.5 N

OpenStudy (mrnood):

@Pomeii00 has it just right for the question as set. (medal) Just for general info - please note that drag is NOT normally proportional to the velocity - it is typically modelled as being proportional to velocity^2 the equation is given as: \[F _{d}= \frac{ C _{d }\rho V ^{2}A }{2 }\] Where Cd is coefficient of drag (depends on shape of body rho is density of medium (e.g. air) V is Velocity A is representative area of body (e.g. cross section) (this is just for reference - the question is fine....)